OpenStudy (anonymous):

calculate the derivative of f if f(x)=sqrt2x+5/x^3+1

OpenStudy (marco26):

right? \[f(x)=\frac{\sqrt{2x+5}}{x^3+1}\]

OpenStudy (marco26):

Use \[\frac{d}{dx}(\frac{u}{v})= \frac{v \frac{d}{dx} (u)-u \frac{d}{dx} (v)}{v^2}\]

OpenStudy (anonymous):

do you just take the derivative of what's inside the square root sign?

OpenStudy (marco26):

no, the whole function

OpenStudy (marco26):

your u will be the square root thing and v the x^3+1